I am building a fence around a rectangular area. The length is twice it's width. The area is 2,450 squared feet. How much fencing materials do I need?
Please show work

Answer by Alan3354(69443)

I am building a fence around a rectangular area. The length is twice it's width. The area is 2,450 squared feet. How much fencing materials do I need?
--------------
It's square feet, not squared feet.
and, it's = it is
--------------------
Area = 2450 = L*W
L = 2W
---------
2450 = 2W*W = 2W^2
W^2 = 1225
W = 35 ft, L = 70 ft
------------
Perimeter = 2L + 2W = 210 feet